A football team won 10 matches out of the total number of matches they played. If their win percentage was 40, then how many matches did they play in all? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the total number of matches played by the team be x.
Given that the team won 10 matches and the winning percentage of the team was 40%.
∴ 40 /100× 𝑥 = 10
⟹ 40x = 10 ×100
⟹ 40x = 1000
⟹ x=1000/40
= 25
∴ The team played 25 matches.

Thanks